OpenStudy (anonymous):

What percent of 26 is 4? 0.26 = n x 4 n = 0.26 x 4 4 = n x 26 26 = 4n

OpenStudy (anonymous):

@phi and @M-e-r-m-a-i-d and @Loser66 and @lila

OpenStudy (anonymous):

@!Star!

OpenStudy (anonymous):

I think it's 4 = n x 26

OpenStudy (anonymous):

@!Star! what do you think?

OpenStudy (anonymous):

I think it is C,

OpenStudy (anonymous):

Okay 3 more please?!?!?!

OpenStudy (anonymous):

K

OpenStudy (anonymous):

What percent of 250 is 70?

OpenStudy (anonymous):

28% right?

OpenStudy (anonymous):

28%

OpenStudy (anonymous):

ok 2 more

OpenStudy (phi):

what percent of 250 is 70 replace "what percent" with "n" n of 250 is 70 replace "of" with times n * 250 is 70 replace is with = n * 250 = 70

OpenStudy (anonymous):

65% of 80 is what number?

OpenStudy (anonymous):

52?

OpenStudy (anonymous):

I have been issued a warning from a someone message me @tatumrocks

OpenStudy (anonymous):

Just message me then You know.

OpenStudy (anonymous):

K